# disposition

### v2

@@ 2.4 Solution concepts in vector optimization
Definition 2.19 (Proper efficiency with respect to a cone-valued map).
Consider a cone-valued map $\mathbb{A}:\Omega\to 2^{Y}$. A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ and $\mathbb{A}(\bar{x})\cap(-K)=\{0\}$. The set of all $\mathbb{A}$-properly efficient solutions of ($P$) is denoted by ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
Remark 2.20.
In Definition 2.19 one can remove the assumption $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ if $\mathbb{A}(\bar{x})$ contains $f[\Omega]-f(\bar{x})$. Notice that ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)={\mathrm{E}ff}(\Omega\mid f,K)$ and ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,\{0\}\cup\operatorname{int}\,K)={\mathrm{W}Eff}(\Omega\mid f,K)$ with $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]-f(\cdot))$ or $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot)).$
Remark 2.21.
For the particular choice of the cone-valued map $\operatorname{\mathbb{A}}:\Omega\to 2^{Y}$, one can see that several well-known concepts of proper efficiency are special cases of Definition 2.19: Hurwicz proper efficiency ([40], 1958), Borwein proper efficiency ([3], 1977), Benson proper efficiency ([1], 1979), extended Borwein proper efficiency ([4], 1980). For more details see, e.g., [5, Sec. 2.4.3], [25, Ch. 3], [26], [49, Sec. 2.4] and references therein.
Beside the proper efficiency concept given in Definition 2.19, we like also to consider a proper efficiency concept in the sense of Henig ([38], 1982). To formulate this concept in the next definition, consider the following family of dilating cones:
@@
Definition 2.22 (Proper efficiency in the sense of Henig).
@@
The following relationships between the different solution concepts are valid. The proof is straightforward.
Proposition 2.23.
Consider a cone $K\subseteq Y$ with $0\notin\operatorname{int}\,K$ (or equivalently, $K\neq Y$). Then, the following assertions hold:
  • $1^{\circ}$.
    $\operatorname{PEff}_{He}(\Omega\mid f,K)\subseteq\operatorname{Eff}(\Omega\mid f,K)\subseteq\operatorname{WEff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    $\operatorname{Eff}(\Omega\mid f,C)=\operatorname{WEff}(\Omega\mid f,K)$ for $C:=(\operatorname{int}\,K)\cup\{0\}$.
  • $3^{\circ}$.
    $\operatorname{PEff}_{He}(\Omega\mid f,K)=\bigcup_{C\in\mathcal{D}(K)}\operatorname{Eff}(\Omega\mid f,C)=\bigcup_{C\in\mathcal{D}(K)}\operatorname{WEff}(\Omega\mid f,C)$.
@@ 3 Scalarization in Vector Optimization
Under the assumptions of Section 2.4, consider the following scalarization of the vector optimization problem ($P$):
@@
where $\varphi:Y\to\mathbb{R}$ is a (frequently parameter-dependent) function. Clearly, to get useful relationships between the problems ($P$) and ($P_{\varphi}$) one needs to impose certain properties on $\varphi$. By solving the scalar problem ($P_{\varphi}$) (with a specific function $\varphi$), one can also get some knowledge about the original vector problem ($P$). The application of such a strategy is called scalarization method, while the function $\varphi$ is called scalarizing function (see e.g. [5], [23], [42], [49]).
In this paper, we like to analyze the relationships between solutions of ($P_{\varphi}$) and the (weakly, properly) efficient solutions of ($P$) for the case that $\varphi$ satisfies certain monotonicity properties. For doing this, we recall monotonicity concepts for the function $\varphi$ (c.f. Jahn [42, Def. 5.1]).
Definition 3.1.
@@
Lemma 3.2.
Consider the subadditive functional $\varphi:Y\to\mathbb{R}$ and let $K,C$ be two cones in $Y$. The following assertions hold:
  • $1^{\circ}$.
    If $\varphi$ is (strictly) $-K$ representing, then $\varphi$ is (strictly) $K$-increasing.
  • $2^{\circ}$.
@@
Proof 3.3.
$1^{\circ}$ We only prove the strict case, since the other case is similar. The case $\operatorname{int}\,K=\emptyset$ is clear. Now assume that $\operatorname{int}\,K\neq\emptyset$. Consider $y^{1},y^{2}\in Y$ with $y^{2}<_{K}y^{1}$, i.e., $y^{2}-y^{1}\in-\operatorname{int}\,K$. Since $\varphi$ is strictly $-K$ representing, this is equivalent to $\varphi(y^{2}-y^{1})<0$. Thus, we derive $\varphi(y^{2})=\varphi(y^{1}+y^{2}-y^{1})\leq\varphi(y^{1})+\varphi(y^{2}-y^{1})<\varphi(y^{1}).$ This shows that $\varphi$ is strictly $K$-increasing. The proofs of $2^{\circ}$ and $3^{\circ}$ are straightforward.
The next scalarization results involving cone-monotone functions are well-known.
Lemma 3.4 (c.f. [50, Lem. 5.6], [27, Lem. 5.2]).
Consider a real-valued function $\varphi:Y\to\mathbb{R}$ and a cone $K\subseteq Y$. Then, the following assertions hold:
@@
    If $\varphi$ is strictly $K$-increasing, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    If $\varphi$ is strongly $K$-increasing, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $3^{\circ}$.
    If $\varphi$ is $K$-increasing, and ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)=\{\overline{x}\}$ for some $\overline{x}\in\Omega$, then $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $4^{\circ}$.
    If $\varphi$ is strictly $C$-increasing or strongly $C$-increasing for some $C\in\mathcal{D}(K)$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
For a (strictly) $-K$ representing function $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$ we have the following result.
Lemma 3.5 (cf. [43, Th. 4.1]).
@@
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)\iff\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
  • $2^{\circ}$.
@@
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)\iff\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
@@
$\begin{cases}\lambda\to\min\\ f(x)\in\lambda k+(a-K),\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$ (3.1)
@@
$\begin{cases}\varphi^{G}_{a-K,k}(f(x))\to\min\\ x\in\Omega,\end{cases}$ (3.2)
where the nonlinear scalarizing function $\varphi^{G}_{a-K,k}:Y\to\overline{\mathbb{R}}:=\mathbb{R}\cup\{-\infty,+\infty\}$ is defined (for $k\in K\setminus\ell(K)$ and $K$ closed) by
$\varphi^{G}_{a-K,k}(y):=\inf\{t\in\mathbb{R}\mid y\in tk+(a-K)\}\quad\text{forall }y\in Y.$ (3.3)
Many facts about the problems (3.1) and (3.2), as well as about the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$, are known (see, e.g., [20, 21], [53], [22], [12, 13], [49], [57], [23]). For the sake of completeness, we recall some important facts below.
Lemma 3.6.
Suppose that $K\subseteq Y$ is a closed, convex cone and take $k\in K\setminus\ell(K)$. Then, the scalar problems (3.1) and (3.2) are equivalent (i.e., each problem has an optimal solution if and only if the other one has an optimal solution, and the optimal solutions $x$ as well as the optimal value of both problems coincide).
Proof 3.7.
Under the given assumptions, the function $\varphi^{G}_{a-K,k}$ has the level set property $f(x)\in\lambda k+(a-K)\iff\varphi^{G}_{a-K,k}(f(x))\leq\lambda$ (see [23, Th. 2.4.1]). Thus, (3.1) is equivalent to
$\begin{cases}\lambda\to\min\\ \varphi^{G}_{a-K,k}(f(x))\leq\lambda,\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$
respectively, to (3.2).
In the next lemma, we state classical relationships between the vector problem ($P$) and the scalar problem (3.1) (see, e.g., Eichfelder [12, Th. 2.1 and Cor. 2.9]).
Lemma 3.8.
@@
  • $1^{\circ}$.
    If $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (3.1) for $a=f(\bar{x})$ and any $k\in K\setminus\ell(K)$. More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1).
  • $2^{\circ}$.
    If $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (3.1) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,K$.
  • $3^{\circ}$.
    Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$. If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $4^{\circ}$.
    If $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$ solves the problem (3.1) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$ and $f(\bar{x})\in\bar{\lambda}k+(a-\operatorname{bd}\,K)$.
Alternatively, one can immediately obtain scalarization results by using the relationships between (3.1) and (3.2), and properties of the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$ defined in (3.3) (see [57, Ch. 6] for an extended overview on this topic).
Lemma 3.9.
Suppose that $K\subseteq Y$ is a closed, convex cone, and take $k\in K\setminus\ell(K)$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$
  • $2^{\circ}$.
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{f(\bar{x})-K,k}(f(x))$.
  • $3^{\circ}$.
    $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the solution set of the problem (3.1) for $a=f(\bar{x})$.
Lemma 3.10.
Suppose that $K\subseteq Y$ is a nontrivial, closed, solid, convex cone, and take $k\in\operatorname{int}\,K$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$
  • $2^{\circ}$.
    $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{f(\bar{x})-K,k}(f(x))$.
  • $3^{\circ}$.
    $(\bar{x},0)$ is a solution of the problem (3.1) for $a=f(\bar{x})$.
The scalar optimization problems of type (3.1) and (3.2) involved in the classical scalarization results given in Lemmata 3.8, 3.9 and 3.10 have the drawback that the parameter $a$ depends on the solution $\bar{x}$ (more precisely, we have $a=f(\bar{x})$). In order to avoid this disadvantage, Jahn [43, Th. 4.2] stated a scalarization result where the parameter $a$ does not depend on the solution $\bar{x}$.
In the following proposition, we present a new scalarization result that contains the result of Jahn [43, Th. 4.2] as a special case (see Remark 3.13 for details).
Proposition 3.11.
Suppose that $K\subseteq Y$ is a convex cone, $\mathcal{P}_{a}\subseteq Y$ is a set of reference points, $\mathcal{P}_{k}\subseteq K\setminus\ell(K)$ is a set of directions and $\mathcal{R}\subseteq\mathbb{R}$ is a set of numbers such that $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of (3.1).
@@
  • $3^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
Proof 3.12.
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is a direct consequence of Lemma 3.8 ($3^{\circ}$).
Now, we show the implication $1^{\circ}\Longrightarrow 2^{\circ}$. Let $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$. Take any $k\in\mathcal{P}_{k}\,(\subseteq K\setminus\ell(K))$. By Lemma 3.8 ($1^{\circ}$), the set of solutions of the problem (3.1) (with $f(\bar{x})$ in the role of $a$) is given by $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$. Note that, for any $x\in\Omega$, the pair ($x,0$) solves (3.1) (with $f(\bar{x})$ in the role of $a$) if and only if ($x,t$) solves (3.1) with $f(\bar{x})-tk$ in the role of $a$ for all $t\in\mathbb{R}$. By our assumption $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, there are $\bar{a}\in\mathcal{P}_{a}$, $\bar{k}\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ such that $f(\bar{x})=\bar{a}+s\bar{k}$. Thus, $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1) with $\bar{a}=f(\bar{x})-s\bar{k}\in\mathcal{P}_{a}$ and $\bar{k}\in\mathcal{P}_{k}$ in the role of $a$ and $k$.
If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 3.6.
Remark 3.13.
@@
For the concept of weak efficiency we have the following result.
Proposition 3.14.
Suppose that $K\subseteq Y$ is a convex cone, $\mathcal{P}_{a}\subseteq Y$ is a set of reference points, $\mathcal{P}_{k}\subseteq\operatorname{int}\,K$ is a set of directions and $\mathcal{R}\subseteq\mathbb{R}$ is a set of numbers such that $f[{\mathrm{W}Eff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $(x,s)$ is a solution of (3.1).

### v3

@@ 2 Preliminaries
Definition 7 (Proper efficiency with respect to a cone-valued map).
Consider a cone-valued map $\mathbb{A}:\Omega\to 2^{Y}$. A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\rm Eff}(\Omega\mid f,K)$ and $\mathbb{A}(\bar{x})\cap(-K)=\{0\}$. The set of all $\mathbb{A}$-properly efficient solutions of ($P$) is denoted by ${\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$.
Remark 4.
In Definition 7 one can remove the assumption $\overline{x}\in{\rm Eff}(\Omega\mid f,K)$ if $\mathbb{A}(\bar{x})$ contains $f[\Omega]-f(\bar{x})$. Notice that ${\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)={\rm Eff}(\Omega\mid f,K)$ and ${\rm PEff}_{\mathbb{A}}(\Omega\mid f,\{0\}\cup\operatorname{int}\,K)={\rm WEff}(\Omega\mid f,K)$ with $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]-f(\cdot))$ or $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot)).$
Remark 5.
For the particular choice of the cone-valued map $\operatorname{\mathbb{A}}:\Omega\to 2^{Y}$, one can see that several well-known concepts of proper efficiency are special cases of Definition 7: Hurwicz proper efficiency ([37], 1958), Borwein proper efficiency ([34], 1977), Benson proper efficiency ([33], 1979), extended Borwein proper efficiency ([35], 1980). For more details see, e.g., [25, Sec. 2.4.3], [31, Ch. 3], [41], [42, Sec. 2.4] and references therein.
Beside the proper efficiency concept given in Definition 7, we like also to consider a proper efficiency concept in the sense of Henig ([36], 1982). To formulate this concept in the next definition, consider the following family of dilating cones:
@@
Definition 8 (Proper efficiency in the sense of Henig).
@@ 3 Scalarization in Vector Optimization
Under the above assumptions, consider the following scalarization of the vector optimization problem ($P$):
@@
where $\varphi:Y\to\mathbb{R}$. Clearly, to get useful relationships between the problems ($P$) and ($P_{\varphi}$) one needs to impose certain properties on $\varphi$. By solving the scalar problem ($P_{\varphi}$) (with a specific function $\varphi$), one can also get certain characterizations for the solutions of the original vector problem ($P$). The application of such a strategy is called scalarization method, while the function $\varphi$ is called scalarizing function (see e.g. [25], [21], [32], [42]).
In this paper, we like to analyze the relationships between solutions of ($P_{\varphi}$) and the (weakly, properly) efficient solutions of ($P$) for the case that $\varphi$ satisfies certain monotonicity properties. For doing this, we recall monotonicity concepts for the function $\varphi$ (c.f. Jahn [32, Def. 5.1]).
Definition 9 ($K$- increasing functional).
@@
The following lemma collects several important relationships between cone monotone functions (based on different cones) as well as cone representing functions. The result will be useful to prove some basic scalarization results based on Bishop-Phelps type cone monotone scalarizing functions in Section 4.1.
Lemma 4.
Consider a function $\varphi:Y\to\mathbb{R}$ and two cones $K,C$ in $Y$. The following assertions hold:
  • $1^{\circ}$
    If $\varphi$ is (strictly) $-K$ representing and subadditive, then $\varphi$ is (strictly) $K$-increasing.
  • $2^{\circ}$
@@
Proof.
$1^{\circ}$ We only prove the strict case, since the other case is similar. The case $\operatorname{int}\,K=\emptyset$ is clear. Now assume that $\operatorname{int}\,K\neq\emptyset$. Consider $y^{1},y^{2}\in Y$ with $y^{2}<_{K}y^{1}$, i.e., $y^{2}-y^{1}\in-\operatorname{int}\,K$. Since $\varphi$ is strictly $-K$ representing, this is equivalent to $\varphi(y^{2}-y^{1})<0$. Thus, we derive $\varphi(y^{2})=\varphi(y^{1}+y^{2}-y^{1})\leq\varphi(y^{1})+\varphi(y^{2}-y^{1})<\varphi(y^{1}).$ This shows that $\varphi$ is strictly $K$-increasing. The proofs of $2^{\circ}$ and $3^{\circ}$ are straightforward. ∎
The next scalarization results involving cone monotone functions (c.f. [48, Lem. 5.6], [52, Lem. 5.2]) are important for deriving the result in Theorem 12.
Lemma 5.
Consider a function $\varphi:Y\to\mathbb{R}$ and a cone $K\subseteq Y$. Then, the following assertions hold:
@@
    If $\varphi$ is strictly $K$-increasing, then ${\rm argmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\rm WEff}(\Omega\mid f,K)$.
  • $2^{\circ}$
    If $\varphi$ is strongly $K$-increasing, then ${\rm argmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\rm Eff}(\Omega\mid f,K)$.
  • $3^{\circ}$
    If $\varphi$ is $K$-increasing, and ${\rm argmin}_{x\in\Omega}\;(\varphi\circ f)(x)=\{\overline{x}\}$ for some $\overline{x}\in\Omega$, then $\overline{x}\in{\rm Eff}(\Omega\mid f,K)$.
  • $4^{\circ}$
    If $\varphi$ is strictly $C$-increasing or strongly $C$-increasing for some $C\in\mathcal{D}(K)$, then ${\rm argmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\rm PEff}_{He}(\Omega\mid f,K)$.
We omit the proof because it is similar to the proofs for [48, Lem. 5.6] and [52, Lem. 5.2], which were established for convex cones in real linear spaces in an algebraic framework.
For a (strictly) $-K$ representing function $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$ we have the following scalarization result which will be used to prove some important characterizations in Propositions 17 and 18.
Lemma 6.
@@
    $\bar{x}\in{\rm Eff}(\Omega\mid f,K)\iff\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
  • $2^{\circ}$
@@
    $\bar{x}\in{\rm WEff}(\Omega\mid f,K)\iff\bar{x}\in{\rm argmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
Remark 6.
Note that Jahn [20, Th. 4.1] stated a similar result as in $1^{\circ}$ of Lemma 6 for the case of a pointed convex cone $K$ in a real linear space $Y$. The proof of Lemma 6 is similar to [20, Th. 4.1].
@@
$\begin{cases}\lambda\to\min\\ f(x)\in\lambda k+(a-K),\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$ (7)
@@
$\begin{cases}\varphi^{G}_{a-K,k}(f(x))\to\min\\ x\in\Omega,\end{cases}$ (8)
where the nonlinear scalarizing function $\varphi^{G}_{a-K,k}:Y\to\overline{\mathbb{R}}:=\mathbb{R}\cup\{-\infty,+\infty\}$ is defined (for $k\in K\setminus\ell(K)$, and $K$ closed and convex) by
$\varphi^{G}_{a-K,k}(y):=\inf\{t\in\mathbb{R}\mid y\in tk+(a-K)\}\quad\text{forall }y\in Y.$ (9)
Many facts about the problems (7) and (8), as well as about the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$, are known (see, e.g., [26, 27], [10], [28], [4, 5], [42], [12], [21]).
Lemma 7.
Suppose that $K\subseteq Y$ is a closed, convex cone and take $k\in K\setminus\ell(K)$. Then, the scalar problems (7) and (8) are equivalent (i.e., each problem has an optimal solution if and only if the other one has an optimal solution, and the optimal solutions $x$ as well as the optimal value of both problems coincide).
Proof.
This fact is well-known (see [12, Th. 5.1.3]). ∎
In the next lemma, we state classical relationships between the vector problem ($P$) and the scalar problem (7) (see, e.g., Eichfelder [4, Th. 2.1 and Cor. 2.9] for the finite dimensional case).
Lemma 8.
@@
  • $1^{\circ}$
    If $\bar{x}\in{\rm Eff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (7) for $a=f(\bar{x})$ and any $k\in K\setminus\ell(K)$. More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (7).
  • $2^{\circ}$
    If $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (7) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,K$.
  • $3^{\circ}$
    Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$. If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (7) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\rm Eff}(\Omega\mid f,K)$.
  • $4^{\circ}$
    If $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$ solves the problem (7) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$ and $f(\bar{x})\in\bar{\lambda}k+(a-\operatorname{bd}\,K)$.
The scalar optimization problems of type (7) and (8) involved in the classical scalarization results (see, e.g., Eichfelder [4, Th. 2.1 and Cor. 2.9]) have the drawback that the parameter $a$ depends on the solution $\bar{x}$ (more precisely, we have $a=f(\bar{x})$). In order to avoid this disadvantage, Jahn [20, Th. 4.2] stated a scalarization result where the parameter $a$ does not depend on the solution $\bar{x}$.
In the following proposition, we present a new scalarization result that contains the result of Jahn [20, Th. 4.2] as a special case (see Remark 7 for details).
Proposition 9.
Suppose that $K\subseteq Y$ is a convex cone, and take sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq K\setminus\ell(K)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\rm Eff}(\Omega\mid f,K)]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$
    $\bar{x}\in{\rm Eff}(\Omega\mid f,K)$.
  • $2^{\circ}$
    There are some $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of (7).
@@
  • $3^{\circ}$
    There are some $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
Proof.
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is a direct consequence of Lemma 8 ($3^{\circ}$).
Now, we show the implication $1^{\circ}\Longrightarrow 2^{\circ}$. Let $\bar{x}\in{\rm Eff}(\Omega\mid f,K)$. Take any $k\in\mathcal{T}\,(\subseteq K\setminus\ell(K))$. By Lemma 8 ($1^{\circ}$), the set of solutions of the problem (7) (with $f(\bar{x})$ in the role of $a$) is given by $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$. Note that, for any $x\in\Omega$, the pair ($x,0$) solves (7) (with $f(\bar{x})$ in the role of $a$) if and only if ($x,t$) solves (7) with $f(\bar{x})-tk$ in the role of $a$ for all $t\in\mathbb{R}$. By our assumption $f[{\rm Eff}(\Omega\mid f,K)]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$, there are $\bar{a}\in\mathcal{P}$, $\bar{k}\in\mathcal{T}$ and $s\in\mathcal{R}$ such that $f(\bar{x})=\bar{a}+s\bar{k}$. Thus, $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (7) with $\bar{a}=f(\bar{x})-s\bar{k}\in\mathcal{P}$ and $\bar{k}\in\mathcal{T}$ in the role of $a$ and $k$.
If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 7. ∎
Remark 7.
@@
For the concept of weak efficiency we have the following new result.
Proposition 10.
Suppose that $K\subseteq Y$ is a convex cone, and take sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq\operatorname{int}\,K$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\rm WEff}(\Omega\mid f,K)]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$
    $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$.
  • $2^{\circ}$
    There are some $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $(\bar{x},s)$ is a solution of (7).